An electro magnetic wave has an electric field given by E=i(225 V/m) sin[(0.077 m^-1)z-(2.3*10^7 rad/s)t]

**********Write down an expression for the magnetic field.
Answer #1

The basic form of these will be

E = E?cos(kx - ?t)

B = B?cos(kx - ?t)

B? = E?/c

E0 = 225

B0= 225/c

So, magnetic field equation is

B=i(225/c V/m) sin[(0.077 m^-1)z-(2.3*10^7 rad/s)t]

= (75*10^-8 T)sin[(0.077 m^-1)z-(2.3*10^7 rad/s)t]
Answer #2

The basic form of these will be

E = E?cos(kx - ?t)

B = B?cos(kx - ?t)

B? = E?/c

E0 = 225

B0= 225/c

So, magnetic field equation is

B=i(225/c V/m) sin[(0.077 m^-1)z-(2.3*10^7 rad/s)t]
Answer #3

c=E/B

B=E/c

=3*10^8/225

=1333333.33 T

So,B=1333333.33sin[(0.077 m^-1)z-(2.3*10^7 rad/s)t]i